You have been tasked to build a rectangular box. The top and bottom are to be built out of 2 cm thick aluminum. The density of aluminum is 2.7 g per cm^3, and the cost $6 per kg. The front and back walls are to be build out of 1 cm thick stainless steel. The density of stainless steel is 7.5 g per cm^3, and the cost $4 per kg. Finally, the 2 remaining walls are to be built out of 1 cm thick copper. The density of copper is 9.0 g per cm^3, and the cost $9 per kg. Find the measurements and total price of the least expensive such box that has volume 1 m^3


1
Expert's answer
2021-10-26T07:06:13-0400

price:

for top and bottom:

"4xy\cdot2.7\cdot0.006=0.0648xy"

for front and back:

"2xz\cdot7.5\cdot0.004=0.06xz"

for 2 remaining walls:

"2yz\cdot9\cdot0.009=0.162yz"


Volume:

"V=xyz=10^6" cm2


"0.0648xy+0.06xz+0.162yz\to min"


surface area:

"A=2xy+2yz+2xz"

"z=V\/(xy)"

"A=2xy+2V\/x+2V\/y"


"\frac{\partial A}{\partial x}=2y-2V/x^2=0\implies y=2V/x^2"


"\frac{\partial A}{\partial y}=2x-2V/y^2=0\implies x=2V/y^2"


"y=2Vy^4\/(4V^2)=y^4\/(2V)"


"y=\sqrt[3]{2V}=\sqrt[3]{2\cdot10^6}=100\sqrt[3]{2}" cm

"x=2\cdot10^6/(10^4\cdot\sqrt[3]{4})=100\sqrt[3]{2}" cm

"z=10^6/(10^4\sqrt[3]{4})=100/\sqrt[3]{4}" cm


Total price:

"P=0.0648xy+0.06xz+0.162yz="

"=0.0648\cdot10^4\sqrt[3]{4}+0.06\cdot 10^4/\sqrt[3]{2}+0.162\cdot 10^4/\sqrt[3]{2}=\$2790.65"